Find the axis of symmetry of the graph of the function.

f(x) = 3x2 + 12x + 7


help
1
Expert's answer
2014-06-05T03:19:42-0400
We can use the next formula to find the axis of symmetry of the function 3x^2+12x+7:
x=-b/2a; We obtain x=-12/(2*3)=-2.


Answer. x=-2.